Find the general solution to the given system:

dx/dt=x+2y

dy/dt=2x+y


1
Expert's answer
2020-12-15T03:33:23-0500

Let us find the general solution to the given system:


"\begin{cases}
\frac{dx}{dt}=x+2y\\
\frac{dy}{dt}=2x+y
\end{cases}"


The first equation of the system is equivalent to "y=\frac{1}{2}(\frac{dx}{dt}-x)". Then "\frac{dy}{dt}=\frac{1}{2}(\frac{d^2x}{dt^2}-\frac{dx}{dt})". Put these in the second equation of the system. We have the following differential equation:


"\frac{1}{2}(\frac{d^2x}{dt^2}-\frac{dx}{dt})=2x+\frac{1}{2}(\frac{dx}{dt}-x)"


which is equivalent to


"\frac{1}{2}\frac{d^2x}{dt^2}-\frac{dx}{dt}-\frac{3}{2}x=0"


and to


"\frac{d^2x}{dt^2}-2\frac{dx}{dt}-3x=0"


Its characteristic equation "k^2-2k-3=0" is equivalent to "(k+1)(k-3)=0", and thus has the solutions "k_1=-1" and "k_2=3."


Therefore, "x(t)=C_1e^{-t}+C_2e^{3t}".


Then "\frac{dx}{dt}=-C_1e^{-t}+3C_2e^{3t}" and


"y=\frac{1}{2}(\frac{dx}{dt}-x)=\frac{1}{2}(-C_1e^{-t}+3C_2e^{3t}-(C_1e^{-t}+C_2e^{3t}))=-C_1e^{-t}+C_2e^{3t}" .


Consequently, the system has the general solution:


"\begin{cases}
x(t)=C_1e^{-t}+C_2e^{3t}\\
y(t)=-C_1e^{-t}+C_2e^{3t}
\end{cases}"